# Capillary waves in miscible fluids

### v1

@@ Abstract
We note that our experiments have been performed in a channel ($W=1$mm) wider than that of [7] ($W=0.25$mm). We therefore start inquiring whether confinement is key for the appearance of waves following the inertial scaling $k\sim\omega^{0}$ or, in other words, whether confinement hampers the capillary regime $k\sim\omega^{2/3}$. We have performed further co-flow experiments employing a narrower channel, $W=$0.25 mm, in the same range of Reynolds numbers ([19]). Notably, we did not observe the onset of a capillary branch in the dispersion relation, consistently with the results of [7] for miscible fluids (Fig. 4-c). For the narrow channel, our data and those of ref. [7] are indeed in excellent quantitative agreement as shown in Fig. 4-c. A further clue of the key role played by confinement is given by the fact that the transition from C to CD waves occurs when the the interface position $Y$ becomes equal to the wave amplitude ([19]), which sharply decreases by approaching further the wall at $y=0$. We conclude that lateral confinement is responsible for suppressing capillary waves under weak tensions and we propose that the mode selection in this case is driven by a minimum-dissipation principle dictating the minimum wavelength emerging for this instability, as it occurs for other instabilities where the capillary stresses at play are very low [31]. It is not surprising therefore that, what have been called inertial regime, defines the upper bound of the unstable region in Fig. 4-a and for the experimental results reported in [7]. Once established the existence of these two kinds of waves in confined co-flows, the CD-regime observed for $Q_{G}\geq$ 8 $\mu$l/l that is characterized by the departure from pure capillary (C) branch towards the inertial one, can be ascribed to the competition between the two modes (C and D). We consider therefore that this is a transition region in a very general stability diagram, as sketched in Fig. 4-d. As a final remark, we emphasize that the same two branches (C and D) delimiting the instability region of the diagram appear in both miscible and immiscible fluids, as shown in Fig. SM7 in [19]. The overlap of the wave dispersions observed in pairs of immiscible and fully miscible fluids support further the importance of interfacial stresses at diffuse interfaces far from equilibrium.
@@
In this letter we reported the existence of waves that emerge spontaneously at the interface between miscible liquids in viscosity-stratified microfluidic flows, following two types of dispersive behavior. Wave propagation is dominated respectively by interfacial stresses for high flow rates and large concentration gradients (C-branch), while dissipation and confinement dominate in the inertial regime, where one single wavenumber emerges. The transition from capillary waves to the inertial regime is generally smooth and occurs when the rate of the fastest fluid is decreased so that diffusion smears out interfaces, interfacial stresses become vanishingly small, and waves get concomitantly more confined. We found very good agreement between the measured values of the effective tension and a model [9] that goes beyond the well-known second order gradient expansion of pressures and mixing free energy in systems far from equilibrium [14, 32]. Our results indicate that interfacial tension in miscible molecular fluids decay much faster right after contact than the temporal decay predicted by square gradient models $\Gamma_{e}\sim t^{-1/2}$, as suggested by other results in static conditions [16], while the square gradient approximation holds at later stages [24]. Capillary waves observed in microfluidic flows of miscible liquids are a novel and unexpected phenomenon, that is put forward as a probe for tensions with unique properties, where steady wave patterns give access to stresses over the millisecond scale after the first liquid-liquid contact, thus paving the way for a thorough understanding of capillary phenomena and flow stability in miscible fluids.
@@ Acknowledgements.
DT warmly thanks Prof. Laura Casanellas for help in manufacturing the microchannels. LC gratefully acknowledges support from the Institut Universitaire de France (IUF). We gratefully acknowledge support from the Centre national d’études spatiales (CNES).

### v3

@@ Abstract
Our experiments have been performed in a channel ($W=1$mm) wider than that of [7] ($W=0.25$mm). We therefore start inquiring whether confinement is key for the appearance of waves following the inertial scaling $k\sim\omega^{0}$, i.e. whether confinement hampers the capillary regime $k\sim\omega^{2/3}$. We have performed further co-flow experiments employing a narrower channel, $W=$0.25 mm, in the same range of Reynolds numbers ([19]). Notably, we did not observe the onset of a capillary branch in the dispersion relation, in excellent qualitative and quantitative agreement with the results of [7] for miscible fluids (Fig. 4-c). A further clue of the key role played by confinement is given by the fact that the transition from C to CD waves occurs as the interface position $Y$ approaches the wave amplitude [19], with the latter sharply decreasing by further approaching the wall at $y=0$. We thus conclude that lateral confinement is responsible for suppressing capillary waves under weak tensions and we propose that the mode selection in this case is driven by a minimum-dissipation principle dictating the minimum wavelength emerging for this instability, as reported for other instabilities where capillary stresses at play are very low [25]. It is not surprising therefore that, what has been called inertial regime, defines the upper (large-$k$) bound of the unstable region in Fig. 4-a and for the experimental results reported in [7]. Having established the existence of capillary and dissipative waves in confined co-flows, the intermediate CD-regime observed for $Q_{G}\geq$ 8 $\mu$l/l can be ascribed to the competition between the C and D modes. We thus consider that the CD modes correspond to a transition region in a very general stability diagram, sketched in Fig. 4-d. As a final remark, we emphasize that the same C and D branches delimiting the instability region of the diagram appear in both miscible and immiscible fluids, see Fig. SM7 in [19]. This overlap of the wave dispersions observed in pairs of immiscible and fully miscible fluids further supports the importance and generality of interfacial stresses at diffuse interfaces far from equilibrium.
@@
To summarize, we have shown that capillary waves, hitherto considered to be a distinctive phenomenon of interfaces between immiscible fluids, emerge also at diffuse boundaries between fully miscible liquids. Their wave propagation is dominated by a transient interfacial tension for high flow rates and large concentration gradients, while dissipation and confinement dominate in the inertial regime, where one single wavenumber emerges and capillary waves are suppressed. The transition from capillary waves to the inertial regime is generally smooth and occurs when the rate of the fastest fluid is decreased so that diffusion smears out interfaces, interfacial stresses become vanishingly small, and waves get concomitantly more confined. We found very good agreement between the measured values of the effective tension and a model [9] that goes beyond the well-known second order gradient expansion of pressures and mixing free energy in systems far from equilibrium [14, 26]. Our results indicate that interfacial tension in miscible molecular fluids decays much faster right after contact than the $t^{-1/2}$ scaling predicted by square gradient models, as hinted at by other results based on the detection of capillary waves in miscible near-critical fluids in static conditions [16], while the square gradient approximation holds at later stages [23]. The capillary wave patterns reported here for microfluidic flows of miscible liquids are a novel and unexpected phenomenon, that we propose as a measurement strategy for accessing transient interfacial tension down to the millisecond time scale. Our work paves the way for a thorough understanding of capillary phenomena and flow stability in miscible fluids, which play a key role in a wide range of research fields of both academic and practical interest, from non-equilibrium thermodynamics [27] and geosciences [28] to oil recovery [29], filtration and flow in porous media (e.g., in a chromatography column [30]), fluid removal [31], and self-propulsion of droplets and colloids of biological interest [32, 33].
@@ Acknowledgements.
DT warmly thanks Prof. Laura Casanellas for help in manufacturing the microchannels. LC gratefully acknowledges support from the Institut Universitaire de France (IUF). We gratefully acknowledge support from the Centre national d’études spatiales (CNES). RG acknowledges support of the Department of Atomic Energy, Government of India, under the project no RTI4001.